OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 A) (1, –2, –3) B) (–4, –1, –2) C)(–3, –2, –3) D)(2, –2, –4)

OpenStudy (anonymous):

Write it as a matrix, and then write that matrix in reduced row echelon form. I'll do it but I'm not an artist so: 1,4,1,-10 3,-3,6,-21 1,2,2,-10 in reduced row echelon form: 1,0,0,-4 0,1,0,-1 0,0,1,-2 x = -4, y = -1, z = -2

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

B) (–4, –1, –2)